Exercise 173

Yappy Company is considering a capital investment
Answer #1

Cash payabck period:

Cumualtive cash flow for year 0 = -320,000

Cumualtive cash flow for year 1 = -320,000 + 62000 = -258,000

Cumualtive cash flow for year 2 = -258,000 + 62,000 = -196,000

Cumualtive cash flow for year 3 = -196,000 + 62,000 = -134,000

Cumualtive cash flow for year 4 = -134,000 + 62,000 = -72,000

Cumualtive cash flow for year 5 = -72,000 + 62,000 = -10,000

Cumualtive cash flow for year 6 = -10,000 + 62,000 = 52,000

10,000 / 62,000 = 0.16

Cash payback period = 5 + 0.16 = 5.16 years

Net present value:

NPV = -320,000 + 62,000 * 5.335

NPV = $10,770

Profitabiltiy index:

Profitability index = Pv of cash flows / initial invesments

Profitability index = 62,000 * 5.335 / 320,000

Profitability index = 330,770 / 320,000

Profitability index = 1.03

Internal rate of return:

Internal rate of return using a financial calculator = 11%

Keys to use in a financial calculator: CF0 = -320,000, CF1 = 62000 F01 = 8, IRR CPT

Annual rate of return

Average net income = (22000 * 8) / 8 = 22,000

Average annual investment = ( 320,000 - 0) / 2 = 160,000

Annual rate of return = 22,000 / 160,000

Annual rate of return = 0.1375 or 13.75%

Investment should be accepted